OpenStudy (shamil98):

\[\int\limits_{0}^{10} \frac{ 1 }{ 3+2 \sin x } dx\] Halp. Can someone give me a step-by-step explanation on how to do this without a graphing calc?

zepdrix (zepdrix):

Hmm I've seen this one a few times. I know it involves some kind of weird substitution like u=tan(x/2) but I'm not sure how to work it. I'm gonna follow the steps on wolfram and see if I can make sense of it. I'm pretty sure harttn would know how to do this one.

OpenStudy (shamil98):

Alright, I guess i'll leave it up for now then.

OpenStudy (kainui):

That's what I tried, I'm playing around with it right now, pretty fun! =)

OpenStudy (anonymous):

@zepdrix mentioned the right substitution. If \(u=\tan\dfrac{x}{2}\), then \(du=\dfrac{1}{2}\sec^2\dfrac{x}{2}~dx\), or \(2\cos^2\dfrac{x}{2}~du=dx\). Now, from this substitution we have the following relations: \[\sin x=\sin\left(2\cdot\frac{x}{2}\right)=2\sin\frac{x}{2}\cos\frac{x}{2}\\ \cos x=\cos\left(2\cdot\frac{x}{2}\right)=\cos^2\frac{x}{2}-\sin^2\frac{x}{2}\] You can draw yourself a triangle to see how you can write these relations in terms of \(u\):  \[\sin\frac{x}{2}=\frac{u}{\sqrt{u^2+1}}\\ \cos\frac{x}{2}=\frac{1}{\sqrt{u^2+1}}\] And so now, \[\sin x=\frac{2u}{u^2+1}\\ \cos x=\frac{1-u^2}{u^2+1}\] Also, you get \(dx=\dfrac{2}{u^2+1}~du\). Now on to the integral: \[\begin{align*}\int\frac{dx}{3+2\sin x}&={\Large\int}\frac{\dfrac{2}{u^2+1}}{3+2\cdot\dfrac{2u}{u^2+1}}du\\ &={\Large\int}\frac{\dfrac{2}{u^2+1}}{\dfrac{3(u^2+1)+4u}{u^2+1}}du\\ &=\int\frac{2}{3(u^2+1)+4u}du\\ &=\int\frac{2}{3u^2+4u+1}du \end{align*}\] The remaining integral can be done by completing the square in the denominator, then applying a trig substitution. There might be an error somewhere in the mix, but the idea is there.